# Dataset Distillation via Knowledge Distillation: Towards Efficient Self-Supervised Pre-training of Deep Networks

### v1

@@ Abstract
Dataset distillation (DD) generates small synthetic datasets that can efficiently train deep networks with a limited amount of memory and compute. Despite the success of DD methods for supervised learning, DD for self-supervised pre-training of deep models has remained unaddressed. Pre-training on unlabeled data is crucial for efficiently generalizing to downstream tasks with limited labeled data. In this work, we propose the first effective DD method for SSL pre-training. First, we show, theoretically and empirically, that naïve application of supervised DD methods to SSL fails, due to the high variance of the SSL gradient. Then, we address this issue by relying on insights from knowledge distillation (KD) literature. Specifically, we train a small student model to match the representations of a larger teacher model trained with SSL. Then, we generate a small synthetic dataset by matching the training trajectories of the student models. As the KD objective has considerably lower variance than SSL, our approach can generate synthetic datasets that can successfully pre-train high-quality encoders. Through extensive experiments, we show that our distilled sets lead to up to 13% higher accuracy than prior work, on a variety of downstream tasks, in the presence of limited labeled data.
@@ 1 Introduction
Dataset distillation (DD) aims to generate a very small set of synthetic images that can simulate training on a large image dataset, with extremely limited memory and compute [37]. This facilitates training models on the edge, speeds up continual learning, and provides strong privacy guarantees [16; 1; 33; 8]. As a result, there has been a surge of interest in developing better DD methods for training neural networks in a supervised manner. However, in many applications, very few labeled example are available. In this case, supervised models often fail to generalize well. Instead, models are pre-trained, using self-supervised learning (SSL), on a large amount of unlabeled data and then adapted to the downstream task using the limited labeled data by training a linear classifier using the labeled examples of each downstream task (linear probe). Remarkably, [2] showed that SSL pre-training, followed by linear probe, can outperform Supervised Learning (SL) by nearly 30% on ImageNet [7] when only 1% of labels are available. More impressively, by only training the linear layer (linear probe), SSL pre-training is able to generalize to a variety of downstream tasks nearly as effectively as full fine-tuning on the downstream task, for a fraction of the cost. Thus, SSL pre-training’s benefits are invaluable in today’s modern ML ecosystem, where unlabeled data is plentiful and it is essential to generalize to a plethora of downstream tasks, effectively and efficiently. With the datasets for SSL being far larger than those for SL, the computational and privacy benefits of dataset distillation for SSL are even more important than they are for SL. Nevertheless, DD for SSL pre-training has remained an open problem.
@@
DD for SSL is, however, very challenging. One needs to ensure that pre-training on the synthetic dataset, distilled from unlabeled data, results in a encoder that yields high-quality representations for a variety of downstream tasks. Existing DD methods for SL generate synthetic data by matching gradients [46; 44; 22] or trajectory [1; 9; 6] of training on distribution of real data [45; 33; 35], or meta-model matching by generating synthetic data such that training on it achieves low loss on real data [37; 25; 23; 48]. Among them, gradient and distribution matching methods heavily rely on labels and will suffer from representation collapse otherwise. Hence, they are not applicable to SSL DD. Very recently, Lee et al. [21] applied meta-model matching to generate synthetic examples for SSL pre-training, and evaluated its performance by fine-tuning the entire pre-trained model on the large labeled downstream datasets. However, we show that SSL pre-training on these distilled sets does not provide any advantage over SSL pre-training on random real examples.
@@
In this work, we address distilling small synthetic datasets for SSL pre-training via trajectory matching. First, we show, theoretically and empirically, that naïve application of trajectory matching to SSL fails, due to the high variance of the gradient of the SSL loss. Then, we rely on insights from knowledge distillation (KD) to considerably reduce the variance of SSL trajectories. KD trains a smaller student network to match the predictions of a larger teacher network trained with supervised learning [12]. In doing so, the student network can match the performance of the larger teacher model.
@@
Finally, we conduct extensive experiments to validate the effectiveness of our proposal MKDT for SSL pre-training. In particular, we distill both low resolution (CIFAR10, CIFAR100) and larger, high resolution datasets (TinyImageNet) down to 2% and 5% of original dataset size and show that, across various downstream tasks, MKDT distilled sets outperform all baselines by up to 13% in the presence of limited labeled data. Moreover, we confirm that the datasets distilled with ConvNets can transfer to other larger architectures such as ResNets. Finally, we demonstrate that MKDT is effective across SSL algorithms (BarlowTwins [43] and SimCLR [29]).
@@ 2.1 Dataset Distillation
There has been a large body of recent work on dataset distillation for supervised learning. These techniques can be broadly characterized into meta-model matching, gradient matching, distribution matching and trajectory matching [32].
@@
Meta-model Matching Meta-model matching generates synthetic data such that a model trained on the synthetic dataset achieves low training loss on the real dataset [37]. The traditional meta-model matching approach is computation and memory inefficient as it requires solving a bi-level optimization problem. Thus, several methods [25; 23; 48] have been proposed to solve the inner-optimization problem in closed form with kernel ridge regression.
@@
Gradient Matching Gradient matching generates synthetic data by matching the gradient of a network trained on the original dataset with the gradient of the network trained on the synthetic dataset [46; 44; 22]. Gradient-matching is done for each class separately, otherwise optimizing the synthetic images to match gradients is not possible [46]. As a result, these methods require labels to be applicable.
@@
Matching Training Trajectories (MTT) MTT, first proposed by Cazenavette et al. [1], generates synthetic data by matching the training trajectories of models trained on the real dataset with that of the synthetic dataset. Cui et al. [6] reduced the memory footprint of MTT and Du et al. [9] minimized the accumulated error in matching trajectories by distilling flatter trajectories.
@@
Distribution Matching Distribution matching generates synthetic data by directly matching the distribution of synthetic dataset and original dataset. One line of work does so by minimizing the maximum mean discrepancy (MMD) between the representations of the synthetic and real data using a large pool of feature extractors [45; 33; 36]. For these methods, distilling examples per class is essential, as without labels, the models trained on the synthetic data suffer from representation collapse and cannot learn any discriminative features [45]. More recently, [40; 47] match the distributions of synthetic and real images using the batch norm statistics of models trained on the full data. While these methods do not distill per class, the distillation loss relies on the labels of the data and is essential to distill data preserving class-discriminative features.
@@
Dataset Distillation for SSL Very recently, KRR-ST [21] applied the kernel-based meta-model matching to distillation for SSL. However, kernel ridge regression, with a relatively unchanging encoder as the kernel function, prevents distilling synthetic data that is useful for training the encoder effectively. We empirically confirm that the encoder learnt by pre-training on these generated examples cannot outperform encoder learnt directly using SSL pre-training on random real images. While KRR-ST also uses a MSE loss to representations instead of directly performing SSL, they claim to do so to mitigate the bias of bi-level optimization in meta-model based matching for SSL. In MKDT, we instead, use the knowledge distillation loss of MSE to representations of a larger teacher model to reduce the high variance of SSL gradients, and thus lower variance trajectories enable trajectory matching.
@@ 2.2 Data-efficient Learning via Subset Selection
Another line of work that enables data-efficient learning is selecting subsets of training data that generalize on par with the full data. This has been extensively studied for supervised learning [5; 34; 28; 24; 39]. At a high level, these works show that difficult-to-learn examples with a higher loss or gradient norm or uncertainty benefit SL the most. More recently, SAS [14] has been proposed for selecting subsets of data for self-supervised contrastive learning (CL). Interestingly, the most beneficial subsets for SL are shown to be least beneficial for self-supervised CL. We use SAS as a baseline and show that the synthetic data distilled by our method can outperform training on these subsets.
@@ 2.3 Knowledge Distillation
Knowledge distillation (KD) is a technique used to transfer knowledge from a large teacher model to a smaller student model, with the aim of retaining high performance with reduced complexity [12]. For supervised learning, some techniques align the student’s outputs with those of the teacher [12], while others concentrate on matching intermediate features [31], attention maps [42] or pairwise distances between examples [26]. Recent works have adapted KD for SSL models [27; 4; 18; 41]. DarkRank [4] approaches KD for SSL as a rank matching problem between teacher and student encoders. PKT [27] and Compress [18] model the similarities in data samples within the representation space as a probability distribution, aiming to align these distributions between the teacher and student encoders. Yu et al. [41] introduced the concept of minimizing the Mean Squared Error (MSE) between the representations of student and teacher encoders. In this work, we rely on KD to enable effective dataset distillation for SSL.
@@ 3 Problem Formulation
Consider a dataset $\mathcal{D}_{\mathsf{real}}=\{\boldsymbol{x}_{i}\}_{i\in[n]}$ of $n$ unlabeled training examples drawn i.i.d. from an unknown distribution. Contrastive SSL methods [43; 2] learn an encoder $f$ that produces semantically meaningful representations by training on $\mathcal{D}_{\mathsf{real}}$. BarlowTwins, in particular, learns these representations using the cross-correlation matrix of the outputs of different augmented views of a given batch of training data:
@@ 4.1 Challenges of Matching SSL Training Trajectories
Matching Training Trajectories for SL MTT [1] generates a synthetic dataset by matching the trajectory of parameters $\hat{\theta}$ of a model trained on the synthetic data with trajectory of parameters $\theta^{*}$ of the model trained on real data (expert trajectory). This matching is guided by the following loss function:
@@
High Variance Gradients Prevent Effective Trajectory Matching for SSL SSL losses rely on interaction between all examples in a batch and consequently have high variance over choices of random batches (c.f., the Barlow-Twins loss in equation 1). As a result, the contribution of examples to the loss and hence their gradients varies significantly based on the rest of examples in the batch [30], unlike SL where each example’s contribution to the loss is independent of other examples. The high variance in gradient over mini-batches in each iteration results in high variance of the trajectories of SSL.
Theoretical Evidence for Higher Variance Gradients in SSL. We now present, in Theorem 4.1, theoretical evidence, in a simplified setting, demonstrating that the variance of the gradient of SSL over mini-batches is indeed greater than that of SL, i.e., $\mathrm{Var}\big{(}\nabla_{W}L_{SSL}(B)\big{)}>\mathrm{Var}\big{(}\nabla_{W}L_{SL}(B)\big{)}$. Proof appears in Appendix C. Appendix D presents a more general version of this analysis, when optimizing with synhronous parallel SGD.
@@
Let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ be a dataset with $n$ examples, where $x_{i}$ is the $i$-th input and $y_{i}\in\{0,1\}$ is its corresponding class label. Assume the data $x_{i}$ are generated using the sparse coding model [38; 15]: for class 0, $x_{i}=e_{0}+\epsilon_{i}$, and for class 1, $x_{i}=e_{1}+\epsilon_{i}$, where $e_{0}$ and $e_{1}$ are basis vectors and $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{N}I)$ is noise. Note that using mean class vectors $e_{0},e_{1}$ w.l.o.g. models the setting of arbitrary mean class vectors that are orthogonal to each other. Each class has $\frac{n}{2}$ examples.
@@
where for a matrix M, $\mathrm{Var}(M):=\mathbb{E}[\left\|M-\mathbb{E}[M]\right\|^{2}]$ as in [11].
Empirical Evidence for Challenges of Matching SSL Trajectories Due to the high variance in the gradient of SSL objectives, the naiv̈e application of MTT to SSL does not succeed. Firstly, the slower convergence caused by high variance gradients necessitates much longer trajectories for both training on real and synthetic data. Secondly, the higher variance of gradients results in greater variance in the weights at the end of trajectories starting from the same initialization (henceforth referred to as variance of trajectories), as confirmed theoretically above. Attempting to match SSL’s longer, higher variance trajectories is extremely challenging, as matching such trajectories results in chaotic updates to the synthetic images. Thus, the synthetic images cannot move away from their initialization meaningfully. Fig. 1(a) shows empirically that the variance of SSL trajectories is larger than that of SL trajectories, across different trajectory lengths. Additionally, the variance of trajectories grows faster, with length of trajectory, for SSL than for SL, exacerbating the problem for longer trajectory matching. Fig. 1(b) compares a simplified distillation using MTT with a single expert trajectory for SSL and SL. Despite extensive hyper-parameter tuning, matching even a single expert trajectory is challenging for SSL, confirmed by the slow decrease of distillation loss. This indicates that the training trajectory on the distilled set is unable to match the training trajectory on the real data for SSL. Fig. 1(c) shows that the difficulty in aligning trajectories is due to the chaotic updates of the synthetic image, as evidenced by the synthetic images being unable to move away from their initialization. To further confirm that the inability to distill effectively is indeed due to the variance of trajectories, we also include a comparison to MTT SSL with 4x larger batch size, which leads to slightly lower variance. Fig. 1(a) confirms that indeed the larger batch size reduces the variance of the trajectories slightly. However, Fig. 1(b) and 1(c) show that reducing the variance of SSL trajectories via larger batch size is insufficient to help distillation since an infeasibly large batch size will likely be required to achieve the necessary low variance trajectories.

### v3

@@ Abstract
Dataset distillation (DD) generates small synthetic datasets that can efficiently train deep networks with a limited amount of memory and compute. Despite the success of DD methods for supervised learning, DD for self-supervised pre-training of deep models has remained unaddressed. Pre-training on unlabeled data is crucial for efficiently generalizing to downstream tasks with limited labeled data. In this work, we propose the first effective DD method for SSL pre-training. First, we show, theoretically and empirically, that naïve application of supervised DD methods to SSL fails, due to the high variance of the SSL gradient. Then, we address this issue by relying on insights from knowledge distillation (KD) literature. Specifically, we train a small student model to match the representations of a larger teacher model trained with SSL. Then, we generate a small synthetic dataset by matching the training trajectories of the student models. As the KD objective has considerably lower variance than SSL, our approach can generate synthetic datasets that can successfully pre-train high-quality encoders. Through extensive experiments, we show that our distilled sets lead to up to 13% higher accuracy than prior work, on a variety of downstream tasks, in the presence of limited labeled data. Code at https://github.com/BigML-CS-UCLA/MKDT
@@ 1 Introduction
Dataset distillation (DD) aims to generate a very small set of synthetic images that can simulate training on a large image dataset, with extremely limited memory and compute (Wang et al., 2018). This facilitates training models on the edge, speeds up continual learning, and provides strong privacy guarantees (Kim et al., 2022; Cazenavette et al., 2022; Sajedi et al., 2023; Dong et al., 2022). As a result, there has been a surge of interest in developing better DD methods for training neural networks in a supervised manner. However, in many applications, very few labeled example are available. In this case, supervised models often fail to generalize well. Instead, models are pre-trained, using self-supervised learning (SSL), on a large amount of unlabeled data and then adapted to the downstream task using the limited labeled data by training a linear classifier using the labeled examples of each downstream task (linear probe). Remarkably, Chen et al. (2020) showed that SSL pre-training, followed by linear probe, can outperform Supervised Learning (SL) by nearly 30% on ImageNet (Deng et al., 2009) when only 1% of labels are available. More impressively, by only training the linear layer (linear probe), SSL pre-training is able to generalize to a variety of downstream tasks nearly as effectively as full fine-tuning on the downstream task, for a fraction of the cost. Thus, SSL pre-training’s benefits are invaluable in today’s modern ML ecosystem, where unlabeled data is plentiful and it is essential to generalize to a plethora of downstream tasks, effectively and efficiently. With the datasets for SSL being far larger than those for SL, the computational and privacy benefits of dataset distillation for SSL are even more important than they are for SL. Nevertheless, DD for SSL pre-training has remained an open problem.
@@
DD for SSL is, however, very challenging. One needs to ensure that pre-training on the synthetic dataset, distilled from unlabeled data, results in a encoder that yields high-quality representations for a variety of downstream tasks. Existing DD methods for SL generate synthetic data by matching gradients (Zhao et al., 2021; Zhao & Bilen, 2021; Lee et al., 2022) or trajectory (Cazenavette et al., 2022; Du et al., 2023; Cui et al., 2022) of training on distribution of real data (Zhao & Bilen, 2023; Sajedi et al., 2023; Wang et al., 2022a), or meta-model matching by generating synthetic data such that training on it achieves low loss on real data (Wang et al., 2018; Nguyen et al., 2021; Loo et al., 2022; Zhou et al., 2022). Among them, gradient and distribution matching methods heavily rely on labels and will suffer from representation collapse otherwise. Hence, they are not applicable to SSL DD. Very recently, Lee et al. (2023) applied meta-model matching to generate synthetic examples for SSL pre-training, and evaluated its performance by fine-tuning the entire pre-trained model on the large labeled downstream datasets. However, we show that SSL pre-training on these distilled sets does not provide any advantage over SSL pre-training on random real examples.
@@
In this work, we address distilling small synthetic datasets for SSL pre-training via trajectory matching. First, we show, theoretically and empirically, that naïve application of trajectory matching to SSL fails, due to the high variance of the gradient of the SSL loss. Then, we rely on insights from knowledge distillation (KD) to considerably reduce the variance of SSL trajectories. KD trains a smaller student network to match the predictions of a larger teacher network trained with supervised learning (Hinton et al., 2015). In doing so, the student network can match the performance of the larger teacher model.
@@
Finally, we conduct extensive experiments to validate the effectiveness of our proposal MKDT for SSL pre-training. In particular, we distill both low resolution (CIFAR10, CIFAR100) and larger, high resolution datasets (TinyImageNet) down to 2% and 5% of original dataset size and show that, across various downstream tasks, MKDT distilled sets outperform all baselines by up to 13% in the presence of limited labeled data. Moreover, we confirm that the datasets distilled with smaller ConvNets can transfer to architectures as large as ResNet-18. Finally, we demonstrate that MKDT is effective across SSL algorithms (BarlowTwins (Zbontar et al., 2021) and SimCLR (Chen et al., 2020)).
@@ 2.1 Dataset Distillation
There has been a large body of recent work on dataset distillation for supervised learning. These techniques can be broadly characterized into meta-model matching, gradient matching, distribution matching and trajectory matching (Sachdeva & McAuley, 2023).
@@
Meta-model Matching Meta-model matching generates synthetic data such that a model trained on the synthetic dataset achieves low training loss on the real dataset (Wang et al., 2018). The traditional meta-model matching approach is computation and memory inefficient as it requires solving a bi-level optimization problem. Thus, several methods (Nguyen et al., 2021; Loo et al., 2022; Zhou et al., 2022) have been proposed to solve the inner-optimization problem in closed form with kernel ridge regression.
@@
Gradient Matching Gradient matching generates synthetic data by matching the gradient of a network trained on the original dataset with the gradient of the network trained on the synthetic dataset (Zhao et al., 2021; Zhao & Bilen, 2021; Lee et al., 2022). Gradient-matching is done for each class separately, otherwise optimizing the synthetic images to match gradients is not possible (Zhao et al., 2021). As a result, these methods require labels to be applicable.
@@
Matching Training Trajectories (MTT) MTT, first proposed by Cazenavette et al. (2022), generates synthetic data by matching the training trajectories of models trained on the real dataset with that of the synthetic dataset. Cui et al. (2022) reduced the memory footprint of MTT and Du et al. (2023) minimized the accumulated error in matching trajectories by distilling flatter trajectories.
@@
Distribution Matching Distribution matching generates synthetic data by directly matching the distribution of synthetic dataset and original dataset. One line of work does so by minimizing the maximum mean discrepancy (MMD) between the representations of the synthetic and real data using a large pool of feature extractors (Zhao & Bilen, 2023; Sajedi et al., 2023; Wang et al., 2022b). For these methods, distilling examples per class is essential, as without labels, the models trained on the synthetic data suffer from representation collapse and cannot learn any discriminative features (Zhao & Bilen, 2023). More recently, (Yin et al., 2024; Zhou et al., 2024; Shao et al., 2024) apply ideas from data-free knowledge distillationLopes et al. (2017) to match the distributions of synthetic and real images using the batch norm statistics of models trained on the full data. While these methods do not distill per class, the distillation loss relies on the labels of the data and is essential to distill data preserving class-discriminative features.
@@
Dataset Distillation for SSL Very recently, KRR-ST (Lee et al., 2023) applied the kernel-based meta-model matching to distillation for SSL. However, kernel ridge regression, with a relatively unchanging encoder as the kernel function, prevents distilling synthetic data that is useful for training the encoder effectively. We empirically confirm that the encoder learnt by pre-training on these generated examples cannot outperform encoder learnt directly using SSL pre-training on random real images. While KRR-ST also uses a MSE loss to representations instead of directly performing SSL, they claim to do so to mitigate the bias of bi-level optimization in meta-model based matching for SSL. In MKDT, we instead, use the knowledge distillation loss of MSE to representations of a larger teacher model to reduce the high variance of SSL gradients, and thus lower variance trajectories enable trajectory matching.
@@ 2.2 Data-efficient Learning via Subset Selection
Another line of work that enables data-efficient learning is selecting subsets of training data that generalize on par with the full data. This has been extensively studied for supervised learning (Coleman et al., 2020; Toneva et al., 2019; Paul et al., 2021; Mirzasoleiman et al., 2020; Yang et al., 2023). At a high level, these works show that difficult-to-learn examples with a higher loss or gradient norm or uncertainty benefit SL the most. More recently, SAS (Joshi & Mirzasoleiman, 2023) has been proposed for selecting subsets of data for self-supervised contrastive learning (CL). Interestingly, the most beneficial subsets for SL are shown to be least beneficial for self-supervised CL. We use SAS as a baseline and show that the synthetic data distilled by our method can outperform training on these subsets.
@@ 2.3 Knowledge Distillation
Knowledge distillation (KD) is a technique used to transfer knowledge from a large teacher model to a smaller student model, with the aim of retaining high performance with reduced complexity (Hinton et al., 2015). For supervised learning, some techniques align the student’s outputs with those of the teacher (Hinton et al., 2015), while others concentrate on matching intermediate features (Romero et al., 2015), attention maps (Zagoruyko & Komodakis, 2017) or pairwise distances between examples (Park et al., 2019). Recent works have adapted KD for SSL models (Passalis & Tefas, 2018; Chen et al., 2017; Koohpayegani et al., 2020; Yu et al., 2019). DarkRank (Chen et al., 2017) approaches KD for SSL as a rank matching problem between teacher and student encoders. PKT (Passalis & Tefas, 2018) and Compress (Koohpayegani et al., 2020) model the similarities in data samples within the representation space as a probability distribution, aiming to align these distributions between the teacher and student encoders. Yu et al. (2019) introduced the concept of minimizing the Mean Squared Error (MSE) between the representations of student and teacher encoders. In this work, we rely on KD to enable effective dataset distillation for SSL.
@@ 3 Problem Formulation
Consider a dataset $\mathcal{D}_{\mathsf{real}}=\{\boldsymbol{x}_{i}\}_{i\in[n]}$ of $n$ unlabeled training examples drawn i.i.d. from an unknown distribution. Contrastive SSL methods (Zbontar et al., 2021; Chen et al., 2020) learn an encoder $f$ that produces semantically meaningful representations by training on $\mathcal{D}_{\mathsf{real}}$. BarlowTwins, in particular, learns these representations using the cross-correlation matrix of the outputs of different augmented views of a given batch of training data:
@@ 4.1 Challenges of Matching SSL Training Trajectories
Matching Training Trajectories for SL MTT Cazenavette et al. (2022) generates a synthetic dataset by matching the trajectory of parameters $\hat{\theta}$ of a model trained on the synthetic data with trajectory of parameters $\theta^{*}$ of the model trained on real data (expert trajectory). This matching is guided by the following loss function:
@@
High Variance Gradients Prevent Effective Trajectory Matching for SSL SSL losses rely on interaction between all examples in a batch and consequently have high variance over choices of random batches (c.f., the Barlow-Twins loss in equation 1). As a result, the contribution of examples to the loss and hence their gradients varies significantly based on the rest of examples in the batch Robinson et al. (2021), unlike SL where each example’s contribution to the loss is independent of other examples. The high variance in gradient over mini-batches in each iteration results in high variance of the trajectories of SSL.
Theoretical Evidence for Higher Variance Gradients in SSL. We now present, in Theorem 4.1, theoretical evidence, in a simplified setting, demonstrating that the variance of the gradient of SSL over mini-batches is indeed greater than that of SL, i.e., $\mathrm{Var}\big{(}\nabla_{W}L_{SSL}(B)\big{)}>\mathrm{Var}\big{(}\nabla_{W}L_{SL}(B)\big{)}$. Proof appears in Appendix C. Appendix D presents a more general version of this analysis, when optimizing with synchronous parallel SGD.
@@
Let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ be a dataset with $n$ examples, where $x_{i}$ is the $i$-th input and $y_{i}\in\{0,1\}$ is its corresponding class label. Assume the data $x_{i}$ are generated using the sparse coding model Xue et al. (2023); Joshi et al. (2024): for class 0, $x_{i}=e_{0}+\epsilon_{i}$, and for class 1, $x_{i}=e_{1}+\epsilon_{i}$, where $e_{0}$ and $e_{1}$ are basis vectors and $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{N}I)$ is noise. Note that using mean class vectors $e_{0},e_{1}$ w.l.o.g. models the setting of arbitrary mean class vectors that are orthogonal to each other. Each class has $\frac{n}{2}$ examples.
@@
where for a matrix M, $\mathrm{Var}(M):=\mathbb{E}[\left\|M-\mathbb{E}[M]\right\|^{2}]$ as in Gower et al. (2020).
Empirical Evidence for Challenges of Matching SSL Trajectories Due to the high variance in the gradient of SSL objectives, the naiv̈e application of MTT to SSL does not succeed. Firstly, the slower convergence caused by high variance gradients necessitates much longer trajectories for both training on real and synthetic data. Secondly, the higher variance of gradients results in greater variance in the weights at the end of trajectories starting from the same initialization (henceforth referred to as variance of trajectories), as illustrated theoretically in the simplified setting above. Attempting to match SSL’s longer, higher variance trajectories is extremely challenging, as matching such trajectories results in chaotic updates to the synthetic images. Thus, the synthetic images cannot move away from their initialization meaningfully. Fig. 1(a) shows empirically that the variance of SSL trajectories is larger than that of SL trajectories, across different trajectory lengths. Additionally, the variance of trajectories grows faster, with length of trajectory, for SSL than for SL, exacerbating the problem for longer trajectory matching. Fig. 1(b) compares a simplified distillation using MTT with a single expert trajectory for SSL and SL. Despite extensive hyper-parameter tuning, matching even a single expert trajectory is challenging for SSL, confirmed by the slow decrease of distillation loss. This indicates that the training trajectory on the distilled set is unable to match the training trajectory on the real data for SSL. Fig. 1(c) shows that the difficulty in aligning trajectories is due to the chaotic updates of the synthetic image, as evidenced by the synthetic images being unable to move away from their initialization. To further confirm that the inability to distill effectively is indeed due to the variance of trajectories, we also include a comparison to MTT SSL with 4x larger batch size, which leads to slightly lower variance. Fig. 1(a) confirms that indeed the larger batch size reduces the variance of the trajectories slightly. However, Fig. 1(b) and 1(c) show that reducing the variance of SSL trajectories via larger batch size is insufficient to help distillation since an infeasibly large batch size will likely be required to achieve the necessary low variance trajectories.